OpenStudy (anonymous):

Write your answer with a positive exponent only. y^(-6)* y^(3)

OpenStudy (mathmale):

\[y^(-6)* y^(3)\rightarrow y ^{-6}y^3=?\]Here both exponential expressions have the same base. What is the rule of exponents that applies to such cases?\[x ^{a}x ^{b}=?\]

OpenStudy (anonymous):

Thank you

OpenStudy (unklerhaukus):

\[x^n\times x^m=x^{n+m}\]